OpenStudy (eric_d):

Find the coefficient of x^4 in the expansion of (2x-1)^15, using binomial theorem

OpenStudy (imstuck):

Hey! Do you know how to use Pascal's triangle?

OpenStudy (eric_d):

Yes..but I need to use the other method

zepdrix (zepdrix):

Using the Binomial Theorem, our Kth term will look like this, \[\Large\rm \left(\begin{matrix}15 \\ k\end{matrix}\right)(2x)^{15-k}(-1)^{k}\] So how do we get a 4th power on our x? What should our k value be so we end up with a power of 4?

OpenStudy (eric_d):

k=11

zepdrix (zepdrix):

\[\Large\rm \left(\begin{matrix}15 \\ 11\end{matrix}\right)(2x)^{15-11}(-1)^{11}\]Ok good.

zepdrix (zepdrix):

\[\Large\rm \left(\begin{matrix}15 \\ 11\end{matrix}\right)(2x)^{4}(-1)^{11}\]

zepdrix (zepdrix):

We've got a ton of numbers floating around, so we'll have to be very careful.

zepdrix (zepdrix):

(-1)^{11} = ?

zepdrix (zepdrix):

Negative to an odd power. What does that simplify to? :d

OpenStudy (eric_d):

negative..

OpenStudy (eric_d):

-1

zepdrix (zepdrix):

Ok good, let's bring it to the front. \[\Large\rm -\left(\begin{matrix}15 \\ 11\end{matrix}\right)(2x)^{4}\]

zepdrix (zepdrix):

How bout the 2x? We have to apply the 4th power to `both` the 2 and the x.

zepdrix (zepdrix):

\[\Large\rm -\left(\begin{matrix}15 \\ 11\end{matrix}\right)2^4x^4\]

zepdrix (zepdrix):

So we'll bring that to the front as well, yes?

zepdrix (zepdrix):

\[\Large\rm -16\left(\begin{matrix}15 \\ 11\end{matrix}\right)x^4\]

zepdrix (zepdrix):

So then we need to deal with this thing: \[\Large\rm \left(\begin{matrix}n \\ k\end{matrix}\right)=\frac{n!}{k!(n-k)!}\]Applying it to our problem:\[\Large\rm \left(\begin{matrix}15 \\ 11\end{matrix}\right)=\frac{15!}{11!(15-11)!}\]

OpenStudy (eric_d):

ok..

zepdrix (zepdrix):

Those factorials are really big nasty numbers. So you'll want to try and cancel things out before expanding out the numbers.

zepdrix (zepdrix):

Here is what I would do....

OpenStudy (eric_d):

what happen to the negative sign

zepdrix (zepdrix):

The -16? We're not looking at that part right now.

zepdrix (zepdrix):

Is that the negative that you're asking about?

OpenStudy (eric_d):

okay

zepdrix (zepdrix):

We're going to rewrite our numerator like this: \[\Large\rm 15!=15\cdot14\cdot13\cdot12\cdot11!\]Does it make sense that they are equal?

OpenStudy (eric_d):

yes coz 15factorial

zepdrix (zepdrix):

\[\Large\rm \left(\begin{matrix}15 \\ 11\end{matrix}\right)=\frac{15\cdot14\cdot13\cdot12\cdot11!}{11!(4)!}\]

zepdrix (zepdrix):

We have a nice cancellation from there,\[\Large\rm \left(\begin{matrix}15 \\ 11\end{matrix}\right)=\frac{15\cdot14\cdot13\cdot12\cdot\cancel{11!}}{\cancel{11!}(4)!}\]

OpenStudy (eric_d):

kk..

zepdrix (zepdrix):

Then maybe just put it into a calculator from there. What value do you get?

OpenStudy (eric_d):

1365

zepdrix (zepdrix):

\[\Large\rm -16\left(\begin{matrix}15 \\ 11\end{matrix}\right)x^4\]Ok great!\[\Large\rm -16\left(1365\right)x^4\]

zepdrix (zepdrix):

and then just multiply to simplify it the rest of the way. Confused at any parts? You were droppin' a lot of dots on me (ok.... ) so I thought maybe you got stuck somewhere lol

OpenStudy (eric_d):

last part..what happens to x^4 -21840x^4?

zepdrix (zepdrix):

Ok great! So you've successfully found the coefficient for x^4 in that expansion.\[\Large\rm -21840\]That's your answer.

OpenStudy (eric_d):

Alrite, thanks for the clear explanation @zepdrix

zepdrix (zepdrix):

np c: